# LETS-C: Leveraging Language Embedding for Time Series Classification

### v1

@@ 5 Conclusion
In this work, we introduced LETS-C, a novel approach that leverages language embeddings for time series classification. To the best of our knowledge, this is the first work to explore the potential of off-the-shelf text embeddings in time series analysis, specifically for classification tasks. By projecting time series data using text embedding models and utilizing a simple yet effective classification head, LETS-C achieves SOTA performance on a well-established benchmark containing 10 datasets across different domains, surpassing 20 baseline methods. Moreover, LETS-C is significantly more lightweight compared to the previous SOTA method, achieving higher accuracy while using much less trainable parameters. Through comprehensive analysis, we demonstrated the effectiveness of LETS-C along different aspects, including its robustness across various text embedding models, the advantage of text embeddings for time series classification, and the trade-off between accuracy and model size. We believe that the findings and insights from this work will inspire further exploration of language embeddings and their potential applications in time series domain, paving the way for the development of more powerful and efficient methods for various time series tasks.
6 Limitations and Future Work
In this work, we adopted the tokenization method proposed by Gruver et al. (2023); however, there may be tokenization strategies better suited specifically for time series data. Future research could investigate these alternatives to optimize performance further (Rajaraman et al., 2024). Additionally, while this study concentrates on using text embeddings exclusively for time series classification, to more comprehensively assess their universality, future research should explore the application of these techniques to a broader spectrum of time series analysis tasks, including forecasting, anomaly detection, and imputation. We expect the societal impact of our work to improve decision-making processes in various sectors, but also practitioners need be careful particularly in sensitive areas like healthcare and personal finance, which raises significant privacy issues. If data is not handled with strict privacy controls, there could be risks of unauthorized access and misuse of personal data.
@@ Appendix A Related Works
Time series classification has been an active research area for decades. Early approaches investigated distance-based approaches Abanda et al. (2019) for time series classification. Some built nearest neighbor classifiers based on explicit time series distance measures such as Dynamic Time Warping (DTW) Wang et al. (2013); Jeong et al. (2011); Berndt & Clifford (1994). Others have used distance kernels instead and learned Support Vector Machines (SVMs) Kampouraki et al. (2008); Bahlmann et al. (2002); Shimodaira et al. (2001), or extracted features and learned linear Dempster et al. (2020) or tree-based classifiers such as eXtreme Gradient Boosting (XGBoost) Chen & Guestrin (2016).
Later, deep learning-based approaches are widely adopted because of their ability to learn complex patterns. Convolutional Neural Networks (CNNs) have proven to be successful in learning local patterns in time series data Wu et al. (2022a); Franceschi et al. (2019); Zhao et al. (2017). Similarly, Multilayer Perceptron (MLP) can provide simple but effective time series classifiers Zhang et al. (2022). Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM) effectively handle long sequence modeling Gu et al. (2021); Lai et al. (2018); Hochreiter & Schmidhuber (1997).
@@ B.5 JapaneseVowels
The Japanese Vowels dataset Kudo et al. , sourced from the UCI Machine Learning Repository, comprises recordings from nine male speakers who pronounced the Japanese vowels ‘a’ and ‘e’. Each utterance was analyzed using a 12-degree linear prediction to extract a 12-dimensional time-series representation, with lengths varying originally from 7 to 29. For consistency, all instances in the dataset have been padded to the maximum length of 29. The objective of the classification task is to identify the speaker; hence each 12-by-29 instance matrix is associated with a single class label, ranging from 1 to 9. This dataset serves as a benchmark for assessing the efficacy of time-series classification models in distinguishing speakers based on LPC cepstrum coefficients obtained from their speech patterns.
@@ B.9 Spoken Arabic Digits
The Spoken Arabic Digits dataset Bedda & Hammami (2010) consists of 8,800 time series data entries derived from the vocal utterances of 88 native Arabic speakers (44 males and 44 females, aged between 18 and 40). Each dataset entry represents one of ten Arabic digits, spoken ten times by each speaker. The dataset captures 13 Mel Frequency Cepstral Coefficients (MFCCs) for each sound snippet, which are extracted under the following audio processing conditions:

### v2

@@ 6 Conclusion
We introduced LETS-C, a novel approach that leverages text embeddings for time series classification. To our knowledge, this is the first exploration of using off-the-shelf text embeddings in time series analysis, particularly for classification. By projecting time series through text embedding models and employing a simple yet effective classification head, LETS-C achieves SOTA performance on a well-established benchmark covering various domains, surpassing 27 baselines. Additionally, LETS-C is significantly more lightweight than previous SOTA methods, achieving higher accuracy with fewer trainable parameters. Our comprehensive analysis highlights LETS-C’s robustness across different text embedding models, the advantage of using text embeddings for time series classification, and the trade-off between accuracy and model size.
7 Limitations and Broader Perspective
Limitations and future work: Our approach has a few key limitations. One is the maximum sequence length constraint of text embedding models, which restricts the handling of longer time series and may necessitate truncation or downsampling, potentially affecting the capture of long-term dependencies. Additionally, while we adhered to established benchmarks for consistency, we did not evaluate our method on a newly constructed or strictly held-out dataset to check for potential data leakage into the OpenAI embeddings. Testing on a novel dataset could further validate our results and ensure that the approach’s effectiveness is not influenced by pre-existing knowledge. Another limitation is the monetary cost associated with using OpenAI’s text embeddings, which we discuss in more detail in Appendix K.
Future research could explore alternative time series tokenization strategies, as well as extensions to other time series tasks. We believe our findings will inspire further exploration of text embeddings in the time series domain, paving the way for more powerful and efficient methods for various time series tasks.
Broader perspective: The embeddings generated may lack interpretability, making it difficult to understand model decisions in high-stakes applications. Additionally, the resources accompanying this study will be responsibly released for research purposes only.
Datasets: The benchmarks used in this study are publicly available from the UEA Archive (Bagnall et al., 2018) and were curated by previous research. Specifically, they include the following datasets: EthanolConcentration (Large et al., 2018), FaceDetection (Rik Henson, 2023), Handwriting (Shokoohi-Yekta et al., 2017), Heartbeat (Liu et al., 2016), JapaneseVowels (Kudo et al., 1999), PEMS-SF (Cuturi, 2011), SelfRegulationSCP1 (Birbaumer et al., 1999), SelfRegulationSCP2 (Birbaumer et al., 1999), SpokenArabicDigits (Bedda and Hammami, 2010), and UWaveGestureLibrary (Liu et al., 2009). We abide by their terms of use.
@@ Appendix A Related Works
Time series classification has been an active research area for decades. Early approaches investigated distance-based approaches Abanda et al. (2019) for time series classification. Some built nearest neighbor classifiers based on explicit time series distance measures such as Dynamic Time Warping (DTW) Wang et al. (2013); Jeong et al. (2011); Berndt and Clifford (1994). Others have used distance kernels instead and learned Support Vector Machines (SVMs) Kampouraki et al. (2008); Bahlmann et al. (2002); Shimodaira et al. (2001), or extracted features and learned linear Dempster et al. (2020) or tree-based classifiers such as eXtreme Gradient Boosting (XGBoost) Chen and Guestrin (2016).
Later, deep learning-based approaches are widely adopted because of their ability to learn complex patterns. Convolutional Neural Networks (CNNs) have proven to be successful in learning local patterns in time series data Wu et al. (2022a); Franceschi et al. (2019); Zhao et al. (2017). Similarly, Multilayer Perceptron (MLP) can provide simple but effective time series classifiers Zhang et al. (2022a). Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM) effectively handle long sequence modeling Gu et al. (2021); Lai et al. (2018); Hochreiter and Schmidhuber (1997).
@@ B.5 JapaneseVowels
The Japanese Vowels dataset Kudo et al. (1999), sourced from the UCI Machine Learning Repository, comprises recordings from nine male speakers who pronounced the Japanese vowels ‘a’ and ‘e’. Each utterance was analyzed using a 12-degree linear prediction to extract a 12-dimensional time-series representation, with lengths varying originally from 7 to 29. For consistency, all instances in the dataset have been padded to the maximum length of 29. The objective of the classification task is to identify the speaker; hence each 12-by-29 instance matrix is associated with a single class label, ranging from 1 to 9. This dataset serves as a benchmark for assessing the efficacy of time-series classification models in distinguishing speakers based on LPC cepstrum coefficients obtained from their speech patterns.
@@ B.9 Spoken Arabic Digits
The Spoken Arabic Digits dataset Bedda and Hammami (2010) consists of 8,800 time series data entries derived from the vocal utterances of 88 native Arabic speakers (44 males and 44 females, aged between 18 and 40). Each dataset entry represents one of ten Arabic digits, spoken ten times by each speaker. The dataset captures 13 Mel Frequency Cepstral Coefficients (MFCCs) for each sound snippet, which are extracted under the following audio processing conditions: